OpenStudy (anonymous):

What is the equation of the line that passes through (5, -2) and (-3, 4)?

OpenStudy (anonymous):

omg thats so easy

OpenStudy (anonymous):

y-y1/y1-y2=x-x1/x1-x2 y1=-2, y2=4, x1=5, x2=-3

OpenStudy (anonymous):

is this the answer? or are these the steps to get the answer?

OpenStudy (anonymous):

oh wait I understand, thank you!